Very short answer question first name any two hydrotropes
 

Grade:7

3 Answers

ROSHAN MUJEEB
3 years ago
Examples ofhydrotropesinclude urea, tosylate, cumenesulfonate and xylenesulfonate. The term hydrotropy was originally put forward by Carl Neuberg to describe the increase in the solubility of a solute by the addition of fairly high concentrations of alkali metal salts of various organic acids.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

When one electron from neutral helium atom is removed, it becomes hydrogen like. The ionisation energy of hydrogen like atom is given by
Eion​ = Z^2Rhc
for helium, Z=2 Rhc=13.6eV
∴ Energy required to liberate second electron from helium
E2​ = Eion ​= 22×13.6eV = −54eV
∴ Energy required to remove both the electrons from helium is
= E1​+E2 ​= (24.6+54.4)eV = 79.0eV

Thanks and Regards
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

When one electron from neutral helium atom is removed, it becomes hydrogen like.
The ionisation energy of hydrogen like atom is given by
Eion​ = Z^2Rhc for helium,
Z=2 Rhc=13.6eV
∴ Energy required to liberate second electron from helium
E2​ = Eion ​= 22×13.6eV = −54eV
∴ Energy required to remove both the electrons from helium is
= E1​+E2
​= (24.6+54.4)eV
= 79.0eV

Thanks and Regards